Question 3 Nashs Trading Post, LLC uses the percentage of receivables basis to record bad debt expense and concludes that 3% of accounts receivable will become uncollectible.Accounts receivable are $439,700 at the end of the year, and the allowance for doubtful accounts has a credit balance of $ 3,041 (a) Prepare the adjusting journal entry to record bad debt expense for the year tb) If the allowance for doubtful accounts had a debit balance of $ 918 instead of a credit balance of $ 3041, prepare the adjusting journal entry for bad debt expense. (Credit account titles are outomatically indented when amount is entered. Do not indent manually) No. Account Titles and Explanation (a) Bad Debt Expense Debit Credit Allowance for Doubtful Accounts (b) Bad Debt Expense Allowance for Doubtful Accounts
Answer #1
Working
Provision = 439700*3% = 13191
Bad debts expense = 13191 - 3041 = 10150
Journal entries
Particulars Debit Credit
a) Bad debts expenses $ 10,150.00
allowance for doubtful accounts $ 10,150.00
(to record Bad debt exp )
b) Bad debts expenses (13191+918) $ 14,109.00
allowance for doubtful accounts $ 14,109.00
( to record bad debts exp )